# TabPFN: One Model to Rule Them All?

### v1

@@ 3.1 Case study I: parameter estimation in a semi-supervised setting
3.2 Case study II: heterogeneous treatment effects estimation
3.2.1 Problem and motivation.
While researchers were previously satisfied with estimating the average effect of a treatment–such as a drug, social intervention, or economic policy–across a population of interest, there has been rising interest in estimating heterogeneous treatment effects (HTE), that is, how different groups of individuals might benefit differentially from the same treatment. Accurate estimation of HTE can allow for personalized treatments, which may lead to better outcomes.
@@ 3.2 Case study II: heterogeneous treatment effects estimation
and the conditional average treatment effect (CATE) function as
$\tau(x)=\mu_{1}(x)-\mu_{0}(x).$
The propensity score is defined as $e(x)=\mathbb{E}(T|X=x)$. In contrast to the average treatment effect (ATE), defined as $\tau=\mathbb{E}[\tau(X)]$, the CATE is able to capture HTE and is hence our estimand of interest.
3.2.2 State-of-the-art.
@@
Foster and Syrgkanis [14] advocated for either the R-Learner or DR-Learner as, unlike the S-, T- and X-Learners, they satisfy Neyman orthogonality, which result in theoretically faster estimation rates so long as the base learners used to estimate the nuisance parameters (e.g. $e(x)$ and $m(x)$) satisfy certain convergence properties. In their numerical experiments, they used Wang et al. [42]’s AutoML algorithm [^5] AutoML uses cross-validation and a smart exploration strategy to select from a collection of the most popular machine learning models and to tune their hyperparameters. as the base learner in the meta-leaner strategies detailed above, and showed that the R-Learner or DR-Learner indeed achieved better performance.
3.2.3 Experimental design.
We repeat Foster and Syrgkanis [14]’s experiments, but add TabPFN as a choice of base learner. We compare the five meta-learner strategies detailed above, together with an “oracle” comprising the R-Learner fitted with the true nuisance functions $m(x)$ and $e(x)$. Note that the optimization problem (3) for the R-Learner is solved via weighted least squares. Since TabPFN currently does not support weighted loss minimization, this means that it cannot be used as a base learner for both the R-Learner and oracle methods. This results in 10 total estimators ($2$ base learners × $6$ methods – $2$ exclusions). Throughout our analysis, we identify each estimator by its method followed by the base learner in parentheses (e.g., ‘S-Learner (TabPFN)’).
In each experiment, we generate a dataset according to the Neyman-Rubin model detailed above, with $X_{i}\sim\text{Unif}([-.5,.5]^{6})$ and $Y_{i}(t)=\mu_{t}(X_{i})+\epsilon_{i}$ for $t=0,1$, where $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$. We explore $6$ different configurations of the functions $e(x)$, $\tau(x)$ and $b(x)=(\mu_{0}(x)+\mu_{1}(x))/2$, but because of space constraints, present only the following two representative cases in the main paper:
@@
Complete specifications of the functional forms and other settings are given in Appendix B. We vary two key experimental parameters: the training set size $n$ and noise level $\sigma^{2}$. For evaluation, we generate a fixed test set of size $10000$ and report the MSE of the estimated treatment effect function evaluated on the test set. The results are averaged over 100 replicates.
@@
3.2.4 Results.
We display the results for Setup A and E in Figure 2 and defer the complete set of experimental results to Table 4 in Appendix B. Most strikingly, across all experimental settings, using TabPFN as the base learner in a meta-learner strategy greatly improves upon its AutoML counterpart. Second, despite its lack of Neyman orthogonality or double robustness, S-Learner (TabPFN) achieves the best accuracy in most settings and is among the top two performers in almost all settings. Our results stand in contrast to those of Foster and Syrgkanis [14] and the rest of the current HTE literature, which has tended to de-emphasize the role played by the base learner because the error it accrues when estimating the nuisance parameters becomes second-order for the R- and DR-Learners under what they consider to be relatively mild conditions. The superior performance of S-Learner (TabPFN) shows that this asymptotic regime is not attained, even under relatively simple simulation settings, and that the choice of base learner can be equally important as the choice of metalearner framework. Indeed, the superior regularization of TabPFN compared to other regression estimators may have a special synergy with the S-Learner strategy, resulting in its superior accuracy, especially in low sample size and signal-to-noise ratio (SNR) settings.
3.3 Case study III: prediction under covariate shift
3.3.1 Problem and motivation.
@@ 3.3 Case study III: prediction under covariate shift
More precisely, we observe a training dataset ${\mathcal{D}}_{s}$ comprising $n$ IID observations from a source distribution $p_{s}(x)p(y|x)$ and wish to predict the responses on a test dataset ${\mathcal{D}}_{t}$ comprising $m$ IID observations from the target distribution $p_{t}(x)p(y|x)$, with accuracy evaluated via prediction MSE. Furthermore, $p_{s}(x)$ and $p_{t}(x)$ are usually unknown and have to be estimated from data.
3.3.2 State-of-the-art.
@@
  • Naive: Fit a gradient boosting regression model on ${\mathcal{D}}_{s}$, tuned using five-fold cross-validation.
  • Importance-Weighted (IW): Fit a gradient boosting regression model on ${\mathcal{D}}_{s}$ similarly to the naive approach, but with the loss function weighted by the true density ratio $p_{t}(x)/p_{s}(x)$.
  • Wang-Oracle: A variant of PL where the true conditional means $\mathbb{E}(Y|X=x)$ are used for $\widetilde{{\mathcal{D}}}_{t}$ instead of imputed values. This serves as an oracle benchmark within Wang [41]’s framework.
@@
$\begin{split}p_{s}(x)=&\frac{5}{6}\mathrm{Unif}(0,0.5)+\frac{1}{6}\mathrm{Unif}(0.5,1),\\ p_{t}(x)=&\frac{1}{6}\mathrm{Unif}(0,0.5)+\frac{5}{6}\mathrm{Unif}(0.5,1).\end{split}$
The response variable follows $Y|X=x\sim{\mathcal{N}}(f(x),1)$, where we investigate five distinct mean functions:
  1. (i)
    $f(x)=\cos(2\pi x)-1$;
  2. (ii)
    $f(x)=\sin(2\pi x)$;
  3. (iii)
    $f(x)=|x-1/2|-1/2$;
  4. (iv)
    $f(x)=f_{1}(x)+f_{2}(x)-2$, $f_{1}(x)=\min\{1,~{}\max\{4x-1,0\}\}$ and $f_{2}(x)=\min\{1,\max\{4x-3,0\}\}$;
  5. (v)
    $f(x)=x\sin(4\pi x)$.
In each experiment, we generate training and test datasets as described above, with the size of the test set fixed at $m=10^{4}$, while we vary the training set size $n\in\{500,600,\dots,1200\}$. The results are averaged over 500 replicates.
3.3.4 Results.
@@
4 TabPFN inductive biases
A classical statistical method makes explicit assumptions about how data is generated, with its prediction performance suffering degradation when those assumptions do not hold. Conversely, when the data contains structure that is not exploited by the method, such as when applying smoothing splines to data generated from a linear model, the prediction performance of the method can be relatively inefficient. Machine learning methods, such as tree-based ensembles and neural networks, are algorithmically defined, which means that these assumptions are implicit rather than explicit. These implicit assumptions are referred to as the method’s inductive bias. Uncovering the inductive biases of ML methods is an active and important area of research, as it tells us when we can expect a given method to do well and when it might perform poorly.
In this section, we attempt to investigate TabPFN’s inductive biases via a collection of numerical experiments. Our takeaways are that
  • In low dimensions, TabPFN seems similar to Gaussian process regression in its ability to adapt flexibly to any functional form, but differs notably in its manner of extrapolation, ease of fitting step functions, and conservativeness at fitting quickly-varying functions.
  • TabPFN can outperform LASSO at high sparsity and low SNR settings by avoiding excessive shrinkage of relevant feature coefficients. On the other hand, it generally performs even worse than LASSO for dense linear regression for $\geq 20$ features.
  • TabPFN is able to break a robustness-efficiency trade-off in classification: It is simultaneously robust to label noise (a type of model misspecification), while being as efficient as linear discriminant analysis (LDA) when it is correctly specified.
4.1 Interpolation and extrapolation
In our first experiment, we examine how TabPFN interpolates and extrapolates from observed data. For simplicity, we demonstrate this in a univariate setting and with respect to four simple functional forms: (i) a linear function, (ii) a quadratic function, (iii) a step function, and (iv) a piecewise linear function. The results for $n=31$ training data points are shown in Figure 4, with the results for $n=11$ and experimental details deferred to Appendix D. For comparison, we also show the results obtained using Gaussian process regression (GPR) with both kernel and noise variance parameters tuned using marginal likelihood. We observe that in all cases, TabPFN fits a smoothly-varying curve that broadly resembles the GPR curve within the interpolation region. The notable differences are:
4.2 Comparison with LASSO regression
In our second experiment, we perform a head-to-head comparison between TabPFN and LASSO when applied to sparse linear regression. The main point is to demonstrate that while TabPFN is seemingly flexible enough to adapt to any type of functional structure, it can nevertheless sometimes recognize and adapt to situations when the dataset contains additional parametric structure, even competing favorably with methods designed specifically to handle that structure. In addition to this, we also observe some other unexpected results.
4.2.1 Experimental design.
@@ 4.2 Comparison with LASSO regression
We broadly follow the experimental design of Hastie, Tibshirani and Tibshirani [18]. We generate $Y=X^{\top}\beta^{*}+\epsilon$, with $\beta^{*}\in{\mathbb{R}}^{p}$ having sparsity $s$ (i.e., $s$ entries are $1$ and the rest are 00). We consider two types of $\beta$:
@@
  • Beta-type I: $\beta^{*}$ has $s$ components equal to $1$, spaced approximately evenly across the $p$ indices, and the rest equal to 00;
  • Beta-type II: $\beta^{*}$ has its first $s$ components equal to $1$, and the rest equal to 00.
@@
4.2.2 Prediction performance comparison.
In Figure 5, we plot the relative test MSE (averaged over $300$ repetitions) of TabPFN to that of LASSO across all experimental settings. We observe that TabPFN generally outperforms LASSO when either the SNR is low or sparsity level is high (i.e. small $s$). However, the performance of TabPFN can be much worse than LASSO for low sparsity (i.e. large $s$) settings, especially when SNR is high. Given that LASSO itself is not especially suited to dense linear regression, this also reveals that TabPFN has a severe inductive bias against dense linear structure. [^6] In other unreported experiments, we also observed that TabPFN performs poorly more generally whenever the regression function depends on a relatively large number of features.
4.2.3 How linear is the TabPFN model?
To further assess and explain these relative performance differences, we wish to inspect the prediction surface of a fitted TabPFN model. To do this, we make use of Accumulated Local Effects (ALE) plots, which summarize the dependence of the model on each individual feature [2]. [^7] ALE plots avoid both the confounding associated with conditional dependence plots and the out-of-sample instability associated with marginal dependence plots. Figure 17 exhibits the results when the features are orthogonal. We observe that for moderate to high SNR ($\geq 1.22$), almost all ALE plots of relevant features are roughly linear. In addition, when $n=500$, the ALE plots for the irrelevant features display much less variation, which suggests that the TabPFN model is indeed well-approximated by a low-dimensional linear surface. For smaller SNR, the ALE plots are generally nonlinear and are not shown.
To further quantify the faithfulness of the linear approximation in the settings identified above ($n=500$, $\text{SNR}\geq 1.22$), we fit an ordinary least squares model to the TabPFN predicted values using only the relevant features. As shown in the Table 2, the resulting test set $R^{2}$ values are at least $0.95$. Furthermore, we can treat the regression coefficients of the linear approximation as estimates of the true regression coefficients. The bias and variance of these estimates are also shown in Table 2. For high to moderate sparsity, we observe that the TabPFN coefficients have smaller bias compared to the LASSO coefficients. Although we report squared bias, we note that all bias values are negative. We thus observe that TabPFN is able to avoid the well-known tendency of LASSO to bias its estimates via shrinkage.
Figure 5: Relative test MSE of TabPFN to LASSO when $n=50$ (left) and $n=500$ (right) under various sparsity and SNR over $1000$ repetitions. Orthogonal features (first row), non-orthogonal feature, beta-type 1 (second row), non-orthogonal feature, beta-type 2 (last row).

### v2

@@ 3.1.4 Results
3.1.5 Real data comparison
To evaluate performance in real-world scenarios, we test the methods on ten OpenML datasets [20], comprising five for regression and five for binary classification. For each dataset, we first establish a ground truth target parameter by fitting a linear model (for regression) or a logistic regression model (for classification) on the full dataset. We then create labeled-unlabeled splits by randomly selecting $n$ observations as labeled data—with $n$ set to 10% of the full dataset for regression and 20% for classification [^4] For numerical stability, we further constrained $n$ to be at least 5,000.—and treating the remainder as unlabeled (using only their features). From these splits, we compute the Vanilla, TabPFN-I, TabPFN-D, and SLZ estimators. In the logistic regression setting, we alternately explore using imputation using the posterior predictive mean for TabPFN-I and TabPFN-D, a procedure that has also been called fractional imputation [47]. [^5] Fractional imputation replaces each observation with a missing binary response by the two pseudo-observations with responses $1$ and 0 and sample weights $P(Y=1\mid X=x)$ and $P(Y=0\mid X=x)$. This yields the same contribution to the score function as imputing the missing response with the single mean value $\sum_{y\in\{0,1\}}y\,P(Y=y\mid X=x)=P(Y=1\mid X=x)$. We repeat this process 50 times and report the average MSE of the estimated coefficients across all repetitions. The data-selection criteria and dataset details are provided in Appendix A.3.
Table 2: Real data SSL: Linear regression.Values represent $\times 10^{-4}$ of the metric
ID Vanilla TabPFN-I TabPFN-D SLZ
189 6.3621 2.3105 1.8860 6.5636
[…]
225 7.0790 1.5684 3.0303 7.0924
1200 1.9185 0.8884 0.7787 1.7529
45052 74.5046 100.3051 64.2511 80.6998
Table 3: Real data SSL: Logistic regression
ID Vanilla TabPFN-I TabPFN-D TabPFN-I TabPFN-D SLZ
mode imputation mean imputation
251 0.0195 2.6087 0.2679 0.0149 0.0181 0.0200
807 0.0036 0.0116 0.0065 0.0020 0.0017 0.0037
816 0.0060 0.1081 0.01550 0.0026 0.0044 0.0060
1046 0.0025 0.0030 0.0022 0.0013 0.0011 0.0025
40983 5.5566 2.3351 2.5356 1.2346 1.9884 5.9513
The results for linear and logistic regression are presented in Table 2 and Table 3, respectively. For linear regression, both TabPFN-I and TabPFN-D typically outperform the Vanilla and SLZ methods. TabPFN-D’s slight advantage over TabPFN-I suggests a benefit from debiasing in these presumably more noisy and complicated settings. For logistic regression, Vanilla outperforms SLZ and TabPFN with mode imputation, whereas TabPFN with mean imputation achieves the best performance. We attribute the weaker performance of mode imputation to its inherent bias: it deterministically assigns a class label even when predictions are uncertain, whereas mean imputation preserves uncertainty by weighting labels according to predicted probabilities, improving stability and accuracy [47].
3.2 Study II: heterogeneous treatment effects estimation
3.2.1 Problem and motivation
While researchers were previously satisfied with estimating the average effect of a treatment–such as a drug, social intervention, or economic policy–across a population of interest, there has been rising interest in estimating heterogeneous treatment effects (HTE), that is, how different groups of individuals might benefit differentially from the same treatment. Accurate HTE estimation enables personalized treatments, which may lead to better outcomes.
@@ 3.2.1 Problem and motivation
and the conditional average treatment effect (CATE) function as $\tau(x)=\mu_{1}(x)-\mu_{0}(x)$. The propensity score is defined as $e(x)=\mathbb{E}(T|X=x)$. Unlike the average treatment effect (ATE) $\tau=\mathbb{E}[\tau(X)]$, the CATE captures HTE and is hence our estimand of interest.
3.2.2 State-of-the-art
@@
Foster and Syrgkanis [21] advocated for either the R-Learner or DR-Learner as, unlike the S-, T- and X-Learners, they satisfy Neyman orthogonality, which result in theoretically faster estimation rates so long as the base learners used to estimate the nuisance parameters (e.g. $e(x)$ and $m(x)$) satisfy certain convergence properties. They used Wang et al. [55]’s AutoML algorithm [^7] AutoML uses cross-validation (CV) and a smart exploration strategy to select from a collection of the most popular machine learning models and to tune their hyperparameters. as the base learner in the meta-leaner strategies detailed above, and showed that the R-Learner or DR-Learner indeed achieved better performance.
3.2.3 Experimental design
We repeat Foster and Syrgkanis [21]’s experiments, but add TabPFN as a choice of base learner. We compare the five meta-learner strategies detailed above, together with an “oracle” comprising the R-Learner fitted with the true nuisance functions $m(x)$ and $e(x)$. Note that the optimization problem (3) for the R-Learner is solved via weighted least squares. Since TabPFN currently does not support weighted loss minimization, this means that it cannot be used as a base learner for both the R-Learner and oracle methods. This results in 10 total estimators ($2$ base learners × $6$ methods – $2$ exclusions). Each estimator is named by its method followed by the base learner in parentheses (e.g., ‘S-Learner (TabPFN)’).
In each experiment, we generate a dataset according to the Neyman-Rubin model detailed above, with $X_{i}\sim\text{Unif}([-.5,.5]^{6})$ and $Y_{i}(t)=\mu_{t}(X_{i})+\epsilon_{i}$ for $t=0,1$, where $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$. We explore $6$ different configurations of the functions $e(x)$, $\tau(x)$ and $b(x)=(\mu_{0}(x)+\mu_{1}(x))/2$, but due to space limits, present only the two representative cases in the main paper:
@@ 3.2.3 Experimental design
Complete specifications of the functional forms and other settings are given in Appendix B. We vary two key experimental parameters: the training set size $n$ and noise level $\sigma^{2}$. For evaluation, we generate a fixed test set of size $10000$ and report the MSE of the estimated treatment effect function evaluated on the test set.
@@
3.2.4 Results
We display the results for Setup A and E in Figure 2 and defer the complete set of experimental results to Table 7 in Appendix B. Most strikingly, across all experimental settings, using TabPFN as the base learner in a meta-learner strategy greatly improves upon its AutoML counterpart. Second, despite its lack of Neyman orthogonality or double robustness, S-Learner (TabPFN) achieves the best accuracy in most settings and is among the top two performers in almost all settings. Our results stand in contrast to those of Foster and Syrgkanis [21] and most of the current HTE literature, which has tended to de-emphasize the role played by the base learner because the error it accrues when estimating the nuisance parameters becomes second-order for the R- and DR-Learners under what they consider to be relatively mild conditions. The superior performance of S-Learner (TabPFN) shows that this asymptotic regime is sometimes not attained even under relatively simple simulation settings, and that the choice of base learner can be as important as the choice of metalearner framework. Indeed, the superior regularization of TabPFN compared to other regression estimators may have a special synergy with the S-Learner strategy, resulting in its superior accuracy, especially in low sample size and signal-to-noise ratio (SNR) settings.
3.2.5 Real data comparison
We evaluate the TabPFN-based meta-learners on the semi-synthetic ACIC 2017 competition benchmark based on the IHDP covariates [8, 17, 26]. The benchmark provides 6,000 datasets generated from real covariates under a controlled set of data-generating processes (DGPs), varying (i) the error structure (IID, non-additive, and group-correlated), (ii) the selection strength $\kappa$ (i.e. variability of the true propensity score), (iii) the treatment effect size, and (iv) the noise level. Each scenario includes ground-truth CATEs, allowing performance assessment by RMSE. We compare the four TabPFN-based meta-learners (S-, T-, X-, and DR-Learner) against the 16 methods described in Hahn, Dorie and Murray [26] that are reported with complete benchmark results. These methods notably include Causal Forests (GRF, Athey, Tibshirani and Wager 3), Targeted Learning (TL, Coyle et al. 14), various versions of BART, and Bayesian Causal Forests (BCF, Hahn, Murray and Carvalho 27), all of which are considered in the literature to be strong baselines. We do not compare with the AutoML-based metal-learners due to the much higher computational overhead required in comparison with using TabPFN.
Figure 3: RMSE of TabPFN-based CATE estimators against SOTA methods on the ACIC 2017 benchmark. Results are shown across panels organized by error type (columns) and selection strength $\kappa$ (rows). Colored bars represent meta-learners built on TabPFN: S-Learner, T-Learner, X-Learner, DR-Learner. The S- , T-, and X-learners enhanced with a propensity score covariate are indicated with the same colors but decorated with a shaded pattern. Bayesian Causal Forests is shown as the shaded yellow bar.
Figure 3 reports the RMSE of each method averaged within each error type and selection strength level. Under weak selection strength, all TabPFN-based estimators match or outperform the baselines. Under strong selection strength, their performance degrades somewhat, which we believe is due to the phenomenon of regularization-induced confounding described in Hahn, Murray and Carvalho [27]—an issue that BCF was specifically designed to fix. Borrowing an idea from BCF, we enhanced the S-, T-, X-Learners by first fitting a propensity score model using a TabPFN classifier and then use the fitted score values as an additional covariate when fitting each base learner model. [^8] The variant is not used for the DR-learner since it already makes use of the propensity score to impute pseudo-outcomes. Under strong selection strength, the resulting methods, especially the X-Learner, all enjoy improved performance that is competitive with the state-of-the-art. Finally, we note that the DR-learner, despite its semi-parametric optimality guarantees, again does not have the best performance in any setting. Full experimental details and per-scenario results are provided in Appendix B.2.
3.3 Study III: prediction under covariate shift
3.3.1 Problem and motivation
@@
More precisely, we observe a training dataset ${\mathcal{D}}_{s}$ comprising $n$ IID observations from a source distribution $p_{s}(x)p(y|x)$ and wish to predict the responses on a test dataset ${\mathcal{D}}_{t}$ comprising $m$ IID observations from the target distribution $p_{t}(x)p(y|x)$, with accuracy evaluated via prediction MSE. Furthermore, $p_{s}(x)$ and $p_{t}(x)$ are usually unknown.
3.3.2 State-of-the-art
@@ 3.3.3 Experimental design
  • Naive: Fit a gradient boosting regression model on ${\mathcal{D}}_{s}$, tuned using five-fold CV.
  • Importance-Weighted (IW): Fit a gradient boosting model on ${\mathcal{D}}_{s}$ similarly to the naive approach, but with the loss function weighted by the true density ratio $p_{t}(x)/p_{s}(x)$.
  • Wang-Oracle: A variant of PL where the true conditional mean $\mathbb{E}(Y|X=x)$ is used for $\widetilde{{\mathcal{D}}}_{t}$ instead of imputed values. This serves as an oracle benchmark within Wang [54].
@@
$\begin{split}p_{s}=&~\frac{5}{6}\mathrm{Unif}(0,0.5)+\frac{1}{6}\mathrm{Unif}(0.5,1),\\ p_{t}=&~\frac{1}{6}\mathrm{Unif}(0,0.5)+\frac{5}{6}\mathrm{Unif}(0.5,1).\end{split}$
The response variable follows $Y|X=x\sim{\mathcal{N}}(f(x),1)$, where we investigate five distinct mean functions: (i) $f(x)=\cos(2\pi x)-1$; (ii) $f(x)=\sin(2\pi x)$; (iii) f(x)=|x1/2|1/2f(x)=|x-1/2|-1/2; (iv) $f(x)=f_{1}(x)+f_{2}(x)-2$, $f_{1}(x)=\min\{1,~\max\{4x-1,0\}\}$ and $f_{2}(x)=\min\{1,\max\{4x-3,0\}\}$; (v) $f(x)=x\sin(4\pi x)$. In each experiment, we generate training and test datasets as described above, with the size of the test set fixed at $m=10^{4}$, while we vary the training set size $n\in\{500,600,\dots,1200\}$. The results are averaged over 500 replicates.
3.3.4 Results
@@
3.3.5 Real data comparison
To assess performance beyond the univariate synthetic setting of Wang [54], we evaluate the methods on two multivariate UCI regression datasets: Airfoil Self-Noise ($n=1503$, 5 covariates) and Concrete Compressive Strength ($n=1030$, 8 covariates). For each dataset, we perform a 70/30 train–test split and resample the test set with replacement using weights $\omega(x)=\exp(x^{\top}\beta)$ following the procedure introduced by Tibshirani et al. [51]. Two shift vectors $\beta$ are considered per dataset, and results are averaged over 10 repetitions. Table 4 reports the mean prediction MSE and standard error across repetitions. Across all datasets and shift settings, TabPFN achieves the best predictive accuracy, often by a substantial margin. Full experimental details are provided in Appendix C.2.
Table 4: Average PMSE (mean $\pm$ standard error) over 10 random splits under covariate shift. Each dataset is evaluated under two $\bm{\beta}$ settings.
Dataset Setting Wang-Oracle PL Naïve IW TabPFN
Airfoil $\bm{\beta}_{1}$ $13.07\pm 2.09$ $14.49\pm 2.15$ $10.25\pm 3.33$ $21.46\pm 7.07$ $\mathbf{3.20\pm 0.84}$
$\bm{\beta}_{2}$ $29\pm 0.72 8.8$$85\pm 0.81$ $3.73\pm 0.57$ $3.46\pm 0.90$ $\mathbf{1.71\pm 0.22}$
Concrete $\bm{\beta}_{1}$ $48.72\pm 15.83$ $109.15\pm 24.72$ $5.50\pm 1.30$ $5.36\pm 1.85$ $\mathbf{2.27\pm 0.45}$
$\bm{\beta}_{2}$ $63.10\pm 19.42$ $116.68\pm 26.49$ $72\pm 0.84 7.7$$78\pm 2.02$ $\mathbf{4.89\pm 0.70}$
4 TabPFN adaptivity
In this section, we illustrate the adaptivity of TabPFN, which we believe provides a compelling explanation for its strong predictive performance across a wide range of tasks, as demonstrated in the previous sections. By adaptivity, we refer to the ability of a procedure to automatically adjust to unknown properties of the underlying data-generating process while still achieving near-optimal performance. This notion has deep roots in nonparametric statistics, where it was classically studied in the context of regression. Foundational work by Lepskii [35], and later by Donoho and Johnstone [16], established estimators that could adapt to unknown smoothness levels of a target function, achieving optimal rates without prior specification of regularity.
However, the form of adaptivity exhibited by TabPFN appears both different and broader. Unlike traditional methods that adapt within a single nonparametric framework—typically modulating local smoothness or bandwidth—TabPFN seems capable of flexibly behaving like a highly nonparametric method when the signal is irregular, yet also exploiting parametric structure when present. In other words, rather than committing to either a parametric or nonparametric regime, TabPFN dynamically adjusts across these regimes, a property that is rare among classical estimators. We will investigate this property in the context of both regression and classification.
4.1 Adaptivity in regression
To investigate adaptivity in regression, we consider the problem of sparse linear regression and compare the performance of TabPFN against that of specialized methods, namely LASSO, SCAD, and MCP regression, across a range of data-generating settings. We find that TabPFN can outperform these methods and can even be used to estimate regression parameters with smaller bias than LASSO. To complement this analysis, we also perform a visual comparison of TabPFN with Gaussian process regression in Section D, observing that the two methods perform similarly, but with subtle differences in their inductive biases.
4.1.1 Experimental design
@@
We follow the experimental design of Hastie, Tibshirani and Tibshirani [28] and generate $Y=X^{\top}\beta^{*}+\epsilon$, with two types of $\beta^{*}\in{\mathbb{R}}^{p}$ having sparsity $s$ (i.e., $s$ entries are $1$ and the rest are 0):
@@
  • Beta-type I: $\beta^{*}$ has $s$ components equal to $1$, spaced approximately evenly across the $p$ indices, and the rest equal to 0;
  • Beta-type II: $\beta^{*}$ has its first $s$ components equal to $1$, and the rest equal to 0.
@@
4.1.2 Prediction performance comparison
In Figure 5, we plot the relative test MSE (averaged over $100$ repetitions) of the comparator methods to that of TabPFN. We focus on the orthogonal feature case, with the results for correlated features deferred to Appendix E.2. Compared to LASSO, we observe that TabPFN tends to perform better for $s\leq 10$, with the performance gap increasing for higher sample size and SNR. Conversely, compared to SCAD and MCP, TabPFN tends to perform better for $s\geq 5$ and for lower sample size and SNR. In other words, TabPFN seems to achieve a type of middle-ground between the two classes of methods.
Figure 5: Relative test MSE compared to TabPFN, using a beta-type I orthogonal design. Rows show results for $n=50$ (top) and $n=500$ (bottom); columns show increasing SNR.